# Distilling the Essence: Efficient Reasoning Distillation via Sequence Truncation

### v1

@@ Abstract
Distilling the reasoning capabilities from a large language model (LLM) to a smaller student model often involves training on substantial amounts of reasoning data. However, distillation over lengthy sequences with prompt (P), chain-of-thought (CoT), and answer (A) segments makes the process computationally expensive. In this work, we investigate how the allocation of supervision across different segments (P, CoT, A) affects student performance. Our analysis shows that selective knowledge distillation over only the CoT tokens can be effective when the prompt and answer information is encompassed by it. Building on this insight, we establish a truncation protocol to quantify computation-quality tradeoffs as a function of sequence length. We observe that training on only the first $50\%$ of tokens of every training sequence can retain, on average, $\approx 94\%$ of full-sequence performance on math benchmarks while reducing training time, memory usage, and FLOPs by about $50\%$ each. These findings suggest that reasoning distillation benefits from prioritizing early reasoning tokens and provides a simple lever for computation-quality tradeoffs. Codes are available at https://github.com/weiruichen01/distilling-the-essence.
@@
Wei-Rui Chen1,2,*, Vignesh Kothapalli2, Ata Fatahibaarzi2, Hejian Sang2, Shao Tang2, Qingquan Song2, Zhipeng Wang2,*, Muhammad Abdul-Mageed 1,*, 1The University of British Columbia, 2LinkedIn weirui.chen@ubc.ca, zhipwang@linkedin.com, muhammad.mageed@ubc.ca
@@ 1 Introduction
Large Language Models (LLMs) benefit from explicit chain-of-thought (CoT) traces to solve multi-step problems (Wei et al., 2023; Kojima et al., 2023). Large Reasoning Models (LRMs) are LLMs optimized to produce such CoT before the answer and achieve superior performance on nontrivial reasoning tasks (Guo et al., 2025), including university-level (Chernyshev et al., 2025) and research-level mathematics (Zhang et al., 2025). However, auto-regressive training and serving of LRMs present significant computational overheads.
Furthermore, combining the reasoning capabilities of such models with test-time scaling to generate long CoT traces, as in the o1 family (OpenAI, 2024), has yielded strong gains on difficult benchmarks (Snell et al., 2024). To this end, distilling reasoning behavior from a powerful teacher LRM into a smaller student model (Hinton et al., 2015; Hsieh et al., 2023; Mukherjee et al., 2023; Mitra et al., 2023; Li et al., 2023) can pave a path for cost-effective deployments in practical settings.
In the long-CoT regime, modern LRMs adopt templates that explicitly separate the prompt, a long CoT segment, and the answer (Guo et al., 2025; Guha et al., 2025; Labs, 2025). Although a growing body of work advocates for the effectiveness of longer CoT traces Chen et al. (2025a) for reasoning tasks (both at test time and at training time), the utility tends to be task dependent and longer chains do not always translate into higher accuracy (Chen et al., 2025b; Sui et al., 2025; Wu et al., 2025). This raises a fundamental yet under-explored question for distillation: ‘which segment (prompt, CoT, answer) is the most effective for supervision, and what fraction of the sequence is actually needed to transfer reasoning abilities efficiently?’
@@
For RQ2, token-budget scaling reveals that using only the first half of tokens retains $\approx 94\%$ of full-sequence performance on average. In contrast, allocating the same budget to the second half of the sequence results in significantly lower accuracy, suggesting that the most critical reasoning signals are concentrated in the early tokens. Collectively, our results position sequence truncation as a new efficiency axis for reasoning distillation, offering a practical path to reduce training time and memory usage. To the best of our knowledge, we are the first to systematically study the supervision of sequence segments and training-time “overthinking” from the perspective of sequence truncation.
@@ 2 Related Work
2.1 Knowledge Distillation and LRMs
Knowledge distillation (KD) is an efficient way to transfer the knowledge of a powerful teacher model to a smaller student model. In the era of LLMs, training and serving full-sized models is extremely compute-intensive. KD mitigates this by enabling a student model to mimic the teacher’s capabilities while requiring substantially fewer resources at both training and inference time. In this sense, KD can be viewed as compressing a large teacher into a smaller model. Researchers have leveraged distillation to improve the reasoning capabilities of LRMs, as a specialized class of LLMs Hsieh et al. (2023); Mukherjee et al. (2023); Mitra et al. (2023).
2.2 Test-time Scaling and Long CoT.
@@ 2.1 Knowledge Distillation and LRMs
A key development in improving LLM reasoning is CoT prompting, which encourages models to generate explicit intermediate reasoning steps before giving an answer (Wei et al., 2023). Such prompting substantially boosts performance on reasoning-intensive tasks in mathematics, science, and code generation (Kojima et al., 2023). Early CoT work typically produced short, highly compressed rationales. The o1 family of models (OpenAI, 2024) demonstrated that allocating substantially more inference-time compute can yield large gains on difficult benchmarks, and subsequent test-time scaling studies (Snell et al., 2024) systematically verified the benefits of generating longer reasoning traces. Following Yeo et al. (2025), we refer to such concise chains as short CoT and to o1-style, slow-thinking approaches as long CoT.
@@
Short CoT methods treat the reasoning trace as an unstructured rationale that is simply concatenated with the answer, using a (prompt, response) template in which both rationale and answer appear in a single segment with no explicit token-level boundary between the two (Wei et al., 2023; Kojima et al., 2023; Wang et al., 2023). More recent works instead split the response into distinct CoT and answer segments, and formulate a (P, CoT, A) template (Guo et al., 2025; Min et al., 2024). For example, DeepSeek-R1 introduces an explicit template in the token stream: the model is trained to place its long CoT inside <think> … </think> and the answer afterward (Guo et al., 2025). STILL-2 (Min et al., 2024), OpenThoughts (Guha et al., 2025), and Bespoke-Stratos (Labs, 2025) adopt similar templates that split and place CoT in <begin_of_thought> … <end_of_thought>.
@@
Despite the empirical success of test-time scaling, several works challenge the idea that “the longer the CoT, the better the performance” and introduce the notion of overthinking at test-time (Chen et al., 2025b; Sui et al., 2025) and at training-time (Wu et al., 2025) when longer CoT does not guarantee better performance. Prior research on LLM reasoning has also noted that not all parts of a CoT are equally valuable. Marjanović et al. (2025) observe that “DeepSeek-R1 often undergoes multiple cycles of self-verification, even when it has already arrived at the correct answer.” This observation suggests that later tokens may carry less useful reasoning signal, and motivate our investigation into whether training on early and partial sequence can match or approach the performance of training on the full sequence.
2.3 Distilling Reasoning Signals.
In the “short CoT” regime, supervision is oftentimes over the response and conditioned on the prompt, as in Orca (Mukherjee et al., 2023), Orca2 (Mitra et al., 2023), Lion (Jiang et al., 2023), and MiniLLM (Gu et al., 2024). With some exceptions, prompts are typically conditioned on but are not included in the loss computation. However, Shi et al. (2024); Huerta-Enochian and Ko (2024) argue that including prompt tokens in the loss is beneficial for downstream performance.
In the “long CoT” setting, supervision can be placed solely on the CoT (Shen et al., 2025), on both CoT and answer  (Li et al., 2025a), or on prompt, CoT, and answer triplet (Xu et al., 2025b, a). However, it is not straightforward to directly compare these approaches, since they differ across experimental factors (datasets, models, hyperparameters, and evaluation standards, among others). Under controlled conditions, we study the roles of supervising different sections (prompt, CoT, and answer) in distilling reasoning models.
@@ 3.1 Formulation
We employ supervised knowledge distillation Hinton et al. (2015); Sanh et al. (2020) [^1] For distillation method choice, see Appendix A.6.2 for details. to transfer the capabilities of a large reasoning teacher model, denoted as $\mathcal{T}$, to a smaller student model that does not undergo any post-training, $\mathcal{S}$. The training objective for the student model is a combination of a cross-entropy loss against the ground-truth labels and a distillation loss that encourages the student to align with the teacher’s output distribution. This composite loss function, $\mathcal{L}$, for a given input sequence $x$ from the training corpus, is formulated as:
@@ 3.2 Experimental Setup
Models. We select the Qwen3 family (Team, 2025b) because its post-trained models are natively trained with two dedicated tokens, <think> and </think>, that explicitly delimit the CoT. In contrast to model families that permit reasoning without explicit start/end markers (e.g., Gemma3 (Team et al., 2025), Mistral (Lin, 2025)), Qwen3 offers a controlled testbed for comparing masking and truncation strategies. We employ two post-trained teachers (Qwen3-32B and Qwen3-8B) and four base student models (Qwen3-8B, 4B, 1.7B, and 0.6B-Base), which are out-of-the-box pre-trained checkpoints without any post-training. To reflect practical compression scenarios, we study seven teacher–student pairs with a parameter ratio of at least $2$ times as shown in Table 1.
@@
Datasets. We distill from two reasoning corpora spanning mathematics, code, and science: OpenThoughts-114k [^3] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k (Openthoughts) (Guha et al., 2025) and Bespoke-Stratos-17k [^4] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k (Bespoke) (Labs, 2025), both produced by DeepSeek-R1. We use these for section-wise supervision, scaling behavior, and early–vs.–late token studies. SkyT1-17k [^5] https://huggingface.co/datasets/NovaSky-AI/Sky-T1_data_17k (SkyT1) (Team, 2025a), produced by QwQ-32B-Preview (Team, 2024), is used exclusively for the early–vs.–late token study (results detailed in Appendix A.4). Each example is a linearized dialogue tree: Prompts (P) are user messages; CoT traces and Answers (A) are assistant messages. System messages are removed. Annotation tags <begin_of_thought> and <end_of_thought> are replaced with <think> and </think> to match the Qwen3 vocabulary. <think> and </think> are used to delimit (prompt, CoT, Answer). Token count statistics of each section for the three datasets are listed in Table 2. Full examples of each dataset and their token distribution of each of the three sections (P, CoT, A) is included in Appendix A.1.
Openthoughts Bespoke SkyT1
Sec Mean Share (%) Mean Share (%) Mean Share (%)
Full 2497.5 - 2165.3 - 2277.1 -
P 116.6 4.7 120.5 5.6 139.1 6.1
CoT 1910.4 76.6 1657.3 76.6 1632.1 71.7
A 468.5 18.8 385.5 17.8 503.9 22.1
@@
Table 2: Token count statistics for three training datasets. “Share (%)” is the average proportion of tokens in each section relative to the full sequence. Distributions are included in Figure 10, 11, and 12 in the Appendix.
@@
To reduce memory usage and ensure that full-length training sequences include all sections (P, CoT, A), we retain only examples with tokenized length below 4k. After filtering, 42,010 examples remain from OpenThoughts, 9,354 from Bespoke, and 9,299 from SkyT1. We then sample 40k, 8k, and 8k instances, respectively, for training, and reserve 1k from each corpus for validation.
@@
Training. Across all settings, we train the student models with knowledge distillation, using Fmchisel framework [^6] https://github.com/linkedin/fmchisel Behdin et al. (2025b, a), for two epochs with AdamW optimizer (learning rate $4\times 10^{-6}$, weight decay $0.05$). We use a batch size of 1, with gradient accumulation, across eight Nvidia H100 (80GB) GPUs [^7] For SkyT1, we use four H100 GPUs with gradient accumulation step being 2. so the effective batch size is $8$. See Appendix A.6.1 for detailed compute resource usage. The checkpoint with the lowest validation loss is selected for evaluation.
@@
Evaluation. We focus our evaluation on mathematical reasoning and consider two widely adopted benchmarks: AIME24888https://huggingface.co/datasets/HuggingFaceH4/aime_2024 MAA (2024) , AIME25999https://huggingface.co/datasets/yentinglin/aime_2025 MAA (2025). To form the baselines, we compare student model performance before and after distillation. During evaluation, we set the maximum generation length to 32,76832{,}768 new tokens, with decoding hyperparameters temperature $0.6$ and top-$p=0.95$, aligning with the recommended values in Qwen3 Huggingface page [^10] https://huggingface.co/Qwen/Qwen3-8B. We report results with the largest sampling budget (64 samples) for both benchmarks to provide a low-variance estimate of the pass@1 metric.

### v2

@@ Abstract
Distilling the capabilities from a large reasoning model (LRM) to a smaller student model often involves training on substantial amounts of reasoning data. However, knowledge distillation (KD) over lengthy sequences with prompt (P), chain-of-thought (CoT), and answer (A) sections makes the process computationally expensive. In this work, we investigate how the allocation of supervision across different sections (P, CoT, A) affects student performance. Our analysis shows that selective KD over only the CoT tokens can be effective when the prompt and answer information is encompassed by it. Building on this insight, we establish a truncation protocol to quantify computation-quality tradeoffs as a function of sequence length. We observe that beyond a specific length, longer training sequences provide marginal returns for downstream performance but require substantially higher memory and FLOPs. To this end, training on only the first $50\%$ of tokens of every training sequence can retain, on average, $\approx 91\%$ of full-sequence performance on math benchmarks while reducing training time, memory usage, and FLOPs by about $50\%$ each. Codes are available at https://github.com/weiruichen01/distilling-the-essence.
@@
Wei-Rui Chen1,2,*,§, Vignesh Kothapalli2, Ata Fatahibaarzi2, Hejian Sang2, Shao Tang2, Qingquan Song2, Zhipeng Wang2,*, Muhammad Abdul-Mageed 1,3,* 1The University of British Columbia, 2LinkedIn, 3Canada Research Chair in NLP and ML weirui.chen@ubc.ca, zhipwang@linkedin.com, muhammad.mageed@ubc.ca
@@
{NoHyper}§§footnotetext: Research conducted during an internship at LinkedIn
@@ 1 Introduction
Distilling the capabilities of LRMs to smaller counterparts has paved a path for cost-effective, practical deployments in recent years (Hinton et al., 2015; Hsieh et al., 2023; Mukherjee et al., 2023; Mitra et al., 2023; Li et al., 2023; Guo et al., 2025). However, datasets curated for this purpose typically comprise long chain-of-thought (CoT) traces (Wei et al., 2023; Kojima et al., 2023) with thousands of tokens and impose significant computational overheads during training. To mitigate such issues, we ask if it is possible to extract meaningful distillation signals from truncated training sequences.
Given an input prompt (P), the CoT trace from a capable LRM reflects the multi-step approach to arrive at an answer (A). Essentially, CoT explicitly showcases the model’s ability to break down a complex task into sub-problems and its self-reflective phases of corrections. Since smaller models are typically incapable of generating such CoT traces Wei et al. (2022), distilling this ability to them can aid in solving reasoning-intensive tasks in resource-constrained settings (Chernyshev et al., 2025; Zhang et al., 2025; OpenAI, 2024; Snell et al., 2024; Guo et al., 2025).
In the long-CoT regime, modern LRMs adopt templates that explicitly separate the prompt, a long CoT segment, and the answer (Guo et al., 2025; Guha et al., 2025; Labs, 2025). Although a growing body of work advocates for the effectiveness of longer CoT traces for reasoning tasks Chen et al. (2025a), their utility tends to be task dependent and longer chains do not always translate into higher accuracy (Chen et al., 2025b; Sui et al., 2025; Wu et al., 2025; Liu and Wang, 2025; Wu et al., 2025). This raises a fundamental yet under-explored question for distillation: ‘which segment (prompt, CoT, answer) is the most effective for supervision, and what fraction of the sequence is sufficient to transfer reasoning abilities efficiently?’
@@
For RQ2, token-budget scaling reveals that beyond a certain length, longer training sequences result in performance plateaus on the downstream math tasks but demand compute resources to scale drastically. In particular, using only the first half of tokens retains $\approx 91\%$ of full-sequence performance on average. In contrast, allocating the same budget to the second half of the sequence results in significantly lower accuracy, suggesting that the most critical reasoning signals are concentrated in the early tokens. Collectively, our results position sequence truncation as a new efficiency axis for reasoning distillation, offering a practical path to reduce training time and memory usage. To the best of our knowledge, we are the first to systematically study the supervision of sequence segments and training-time “overthinking” from the perspective of sequence truncation.
@@ 2 Related Work
2.1 Long Chain-of-Thought
@@
A key development in improving Large Language Model (LLM) reasoning is CoT prompting, which encourages models to generate explicit intermediate reasoning steps before giving an answer (Wei et al., 2023). Such prompting substantially boosts performance on reasoning-intensive tasks in mathematics, science, and code generation (Kojima et al., 2023). Early CoT work typically produced short, highly compressed rationales. The o1 family of models (OpenAI, 2024) demonstrated that allocating substantially more inference-time compute can yield large gains on difficult benchmarks, and subsequent test-time scaling studies (Snell et al., 2024) systematically verified such benefits. Following Yeo et al. (2025), we refer to such concise chains as short CoT and to o1-style, slow-thinking approaches as long CoT.
@@
Short CoT typically concatenates the reasoning trace with the answer, using a (P, CoT + A) template in which both rationale and answer appear in a single segment with no explicit token-level boundary (Wei et al., 2023; Kojima et al., 2023; Wang et al., 2023). More recent works instead split the response into distinct CoT and answer segments, and formulate a (P, CoT, A) template (Guo et al., 2025; Min et al., 2024). For example, DeepSeek-R1 introduces an explicit template in the token stream: the model is trained to place its long CoT inside <think> … </think> and the answer afterward (Guo et al., 2025). STILL-2 (Min et al., 2024), OpenThoughts (Guha et al., 2025), and Bespoke-Stratos (Labs, 2025) adopt similar templates that split and place CoT in <begin_of_thought> … <end_of_thought>. Such separation allows long reasoning steps to be included in the CoT segment, whereas a concise final solution is placed in the answer (A) segment. Beyond the common practice of utilizing the full sequence (P, CoT, A) for distilling reasoning capabilities to smaller models Guo et al. (2025), our work takes a systematic approach to isolating the utility of each of these segments.
2.2 Distilling Reasoning Signals
In the era of LRMs, training and serving full-sized models is extremely compute-intensive. To mitigate such bottlenecks, Knowledge Distillation (KD) Hinton et al. (2015) offers an efficient way to transfer the knowledge of a powerful teacher LRM to a smaller student model Hsieh et al. (2023); Mukherjee et al. (2023); Mitra et al. (2023). In the short CoT regime, supervision from the teacher model is oftentimes emphasized over the response and conditioned on the prompt, as in Orca (Mukherjee et al., 2023), Orca2 (Mitra et al., 2023), Lion (Jiang et al., 2023), and MiniLLM (Gu et al., 2024). With some exceptions, prompts are typically conditioned on but are not included in the loss computation.
@@ 2.2 Distilling Reasoning Signals
In the long CoT setting, there are two main aspects to consider. The choice of segments for supervision and the choice of reasoning datasets for distillation. For the first aspect, supervision (i.e, loss calculation) can be placed solely on the CoT segment (Shen et al., 2025), on both CoT and answer segments  (Li et al., 2025b), or on the entire sequence (prompt, CoT, and answer triplet) (Xu et al., 2025b, a). For the second aspect of choosing datasets, we note that several works have challenged the idea that “longer CoT implies better performance” and introduced the notion of reasoning with overthinking (Chen et al., 2025b; Sui et al., 2025; Liu and Wang, 2025; marjanović2025deepseekr1thoughtologyletsthink). Utilizing sequences that contain such overthinking signals for distillation can be detrimental to model performance (Wu et al., 2025). Prior research on LLM reasoning has also noted that not all parts of a CoT are equally valuable. marjanović2025deepseekr1thoughtologyletsthink observe that “DeepSeek-R1 often undergoes multiple cycles of self-verification, even when it has already arrived at the correct answer.” This observation suggests that later tokens may not carry significantly useful reasoning signals. These observations motivate our controlled investigation into whether distillation of early and partial sequences can match or approach the efficacy of full-length sequences. In particular, we focus on computation-quality tradeoffs with increasingly long sequences.
@@ 3.1 Formulation
We employ supervised knowledge distillation Hinton et al. (2015); Sanh et al. (2020) [^1] For method choice, see Appendix A.4 and A.5.3 for details. to transfer the capabilities of a large reasoning teacher model, denoted as $\mathcal{T}$, to a smaller student model that did not undergo any post-training, $\mathcal{S}$. The training objective for the student model is a combination of a cross-entropy loss against the ground-truth labels and a distillation loss that encourages the student to align with the teacher’s output distribution. This composite loss function, $\mathcal{L}$, for a given input sequence $x$ from the training corpus, is given by:
@@ 3.2 Experimental Setup
Models. We select the Qwen3 family (Team, 2025b) because its post-trained models are natively trained with two dedicated tokens, <think> and </think>, that explicitly delimit the CoT. In contrast to model families that permit reasoning without explicit start/end markers, such as Gemma3 (Team et al., 2025) and Mistral (Lin, 2025), Qwen3 offers a controlled testbed for comparing masking and truncation strategies. We employ two post-trained teachers (Qwen3-32B and Qwen3-8B) and four base student models (Qwen3-8B, 4B, 1.7B, and 0.6B-Base), which are out-of-the-box pre-trained checkpoints without any post-training. To reflect practical distillation scenarios, we consider the seven teacher-student pairs as shown in Table 1.
@@
Datasets. We distill from five reasoning corpora: OpenThoughts-114k [^3] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k (Openthoughts) (Guha et al., 2025), Bespoke-Stratos-17k [^4] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k (Bespoke) (Labs, 2025), Synthetic-1555https://huggingface.co/datasets/PrimeIntellect/SYNTHETIC-1 (Mattern et al., 2025), Llama-Nemotron-Post-Training-Dataset [^6] https://huggingface.co/datasets/nvidia/Llama-Nemotron-Post-Training-Dataset (Nemotron) (Bercovich et al., 2025) and SkyT1-17k [^7] https://huggingface.co/datasets/NovaSky-AI/Sky-T1_data_17k (SkyT1) (Team, 2025a). The first four datasets were generated by DeepSeek-R1, whereas the last was produced by QwQ-32B-Preview (Team, 2024). We utilize Openthoughts and Bespoke for the full suite of experiments on sequence truncation and segment utility. To examine the generalizability of our findings, we additionally employ Synthetic-1, Nemotron, and SkyT1 exclusively for such analysis. Each sequence/example is a linearized dialogue tree: Prompts (P) are user messages; CoT traces and Answers (A) are assistant messages. System messages are removed. Annotation tags <begin_of_thought> and <end_of_thought> are replaced with <think> and </think> to match the Qwen3 vocabulary. <think> and </think> are used to delimit (prompt, CoT, Answer). Token count statistics of each section for the five datasets are listed in Table 2. Full examples of each dataset and their token distribution of each of the three sections (P, CoT, A) are included in Appendix A.1.
Openthoughts Bespoke Synthetic-1 Nemotron SkyT1
Sec Mean % Mean % Mean % Mean % Mean %
Full 2497.5 - 2165.3 - 1854.0 - 2437.8 - 2277.1 -
P 116.6 4.7 120.5 5.6 100.2 5.4 87.3 3.6 139.1 6.1
CoT 1910.4 76.6 1657.3 76.6 1402.8 75.7 2073.3 85.1 1632.1 71.7
A 468.5 18.8 385.5 17.8 349.1 18.8 275.2 11.3 503.9 22.1
@@
Table 2: Token count statistics for five training datasets. “%” is the average proportion of tokens in each section relative to the full sequence. Distributions are included in Appendix A.1.2.
@@
To reduce memory usage and ensure that full-length training sequences include all sections (P, CoT, A), we retain only examples with tokenized length below 4k. After filtering, we then sample 40k, 8k, 16k, 16k, and 8k instances from Openthoughts, Bespoke, Synthetic-1, Nemotron, and SkyT1, respectively, for training, and reserve 1k from each corpus for validation.
@@
Training. Across all settings, we train the student models with knowledge distillation, using Fmchisel framework [^8] https://github.com/linkedin/fmchisel Behdin et al. (2025b, a), for two epochs with AdamW optimizer (learning rate $4\times 10^{-6}$, weight decay $0.05$). We use a batch size of 1, with gradient accumulation, across eight Nvidia H100 (80GB) GPUs so the effective batch size is $8$. See Appendix A.5.2 for detailed compute resource usage. The checkpoint with the lowest validation loss is selected for evaluation.
@@
Evaluation. We focus our evaluation on mathematical reasoning and consider two widely adopted benchmarks: AIME24999https://huggingface.co/datasets/HuggingFaceH4/aime_2024 MAA (2024) and AIME25101010https://huggingface.co/datasets/yentinglin/aime_2025 MAA (2025). To form the baselines, we compare student model performance before and after distillation. During evaluation, we set the maximum generation length to 32,76832{,}768 new tokens, with decoding hyperparameters temperature $0.6$ and top-$p=0.95$, aligning with the recommended values in Qwen3 Huggingface page [^11] https://huggingface.co/Qwen/Qwen3-8B. We report results with the largest sampling budget (64 samples) for both benchmarks using lighteval [^12] https://github.com/huggingface/lighteval.